# language model LLM large language model

### v1

@@ Abstract
Recent studies show that large language models (LLMs) struggle with technical standards in telecommunications. We propose a fine-tuned retrieval-augmented generation (RAG) system based on the Phi-2 small language model (SLM) to serve as an oracle for communication networks. Our developed system leverages forward-looking semantic chunking to adaptively determine parsing breakpoints based on embedding similarity, enabling effective processing of diverse document formats. To handle the challenge of multiple similar contexts in technical standards, we employ a re-ranking algorithm to prioritize the most relevant retrieved chunks. Recognizing the limitations of Phi-2’s small context window, we implement a recent technique, namely SelfExtend, to expand the context window during inference, which not only boosts the performance but also can accommodate a wider range of user queries and design requirements from customers to specialized technicians. For fine-tuning, we utilize the low-rank adaptation (LoRA) technique to enhance computational efficiency during training and enable effective fine-tuning on small datasets. Our comprehensive experiments demonstrate substantial improvements over existing question-answering approaches in the telecom domain, achieving performance that exceeds larger language models such as GPT-4 (which is about 880 times larger in size). This work presents a novel approach to leveraging SLMs for communication networks, offering a balance of efficiency and performance. This work can serve as a foundation towards agentic language models for networks.
@@ I Introduction
Recent studies indicate that while state-of-the-art LLMs perform well on general telecommunications queries, they struggle with questions related to technical standards in the field [11, 12]. We believe this is mainly due to the fact that standard-type knowledge and system specifications do not exist in common research papers and other publications, which serve as the main learning sources for LLMs. For instance, with respect to telecom systems, the polysemy of abbreviations (e.g., SAP: service access sample vs. system application and protocol) can hinder the model from inferring a correct answer. Additionally, the LMs training on generalized knowledge can interfere with its performance in specialized domain tasks. For instance, certain protocols and methods in telecom-specific domains do not necessarily follow the generally followed-upon practices in broader contexts. A model trained on generalized knowledge may not adequately capture these domain-specific nuances and practices.
We propose a carefully developed Phi-2 based fine-tuned RAG system to serve as an oracle for communication networks. To the best of our knowledge, this is the first work to present a fine-tuned RAG system for communication networks. Previous works focus on presenting a frozen RAG framework or fine-tuning an LM. Our RAG system leverages a forward-looking semantic chunking (or parsing) strategy that adaptively determines breakpoints between sentences based on embedding similarity. This approach enables the system to effectively process documents with diverse formatting. In the Third Generation Partnership Project (3GPP) documents, a query can often relate to multiple similar contexts, as discussions and paragraphs on related topics may appear in various sections or be phrased similarly. Therefore, we utilize a re-ranking algorithm to further rank the retrieved chunks based on their relevance to the input query.
@@ II Related Works
In [11], the introduction of the TeleQnA dataset marks a significant advancement in evaluating QnA tasks for telecommunications. The TeleQnA dataset contains 10,000 MCQs about telecommunication systems, curated from the various sources such as 3GPP and research papers. The dataset is verified by telecom human-in-the-loop experts. Another dataset, namely TSpec-LLM, is recently released [16]. The authors develop an automated framework to generate QnA pairs from 3GPP specifications, then test a naive RAG architecture to assess the quality of their dataset. In [17], Gajjar et al. introduce ORAN-Bench-13K, a dataset dedicated for the evaluation of open radio access networks (O-RAN) tasks. The dataset is based on 116 O-RAN specification documents and contains 13,000 pairs of MCQs, based on which ORANSight is developed, a RAG-based framework.
@@ III-E Extending the Context Window with SelfExtend
However, SLMs typically struggle to generalize effectively to input sequences longer than those encountered during training. This presents as a challenge during inference with long contexts. The context window of SLMs are often short. For example, Phi-2 has a context window of 2048 tokens. Semantic chunking does not guarantee a fixed chunk size, and therefore, might exceed the context window of the SLMs.
@@ III-F The Generator: Fine-tuned Phi-2 with Multiple Contexts
A substantial portion of telecom key terms and special language is confined to specification documents and white papers, neither of which LMs are heavily trained on. We fine-tune Phi-2 to overcome this and to adapt the model to recognize telecom terminology.
@@
The concept underlying LoRA is that pre-trained LMs possess a low ‘intrinsic dimension’ [26]. That is, the model’s essential information is concentrated in a smaller subspace, even if the overall parameter space is high-dimensional. LoRA harnesses this observation and focus the model updates on a smaller only a subset of the learning parameters. During fine-tuning, weight updates $W_{new}$ are represented as
@@ III-G Prompt Engineering
Prompt engineering is crucial for optimizing the performance of LLMs like Phi-2 in telecom applications. In this work, we supplement the inputted question with chunks of relevant context, and a set of instructions for the SLMs to follow, forming the prompt. n exemplary prompt is outlined in 3. This approach helps unify the format of the SLM’s output, focus its attention on the appearance of critical terms, and encourage it to reply on the context rather than its prior knowledge. This prompt results in focused outputs, ensuring accurate and relevant responses for complex telecom inquiries.
@@ IV-A Settings
The RAG framework relies on around 550 3GPP documents up to Release 18. For fine-tuning, we utilize the TeleQnA dataset which contains 10,000 multiple-choice questions [11]. The testing set features another 2,000 multiple-choice questions that focus on 3GPP standards. For fine-tuning, we set the following parameters: weight decay of 0.01, batch size of 32, dropout rate of 0.05, and learning rate of $10^{-4}$. For LoRA, we set the rank to 32 and the alpha to 64. For semantic chunking, there are two hyper-parameters, namely the breakpoint percentile threshold and the buffer size. The former represents the percentile of cosine dissimilarity that must be exceeded between a group of sentences and the next to form a node. The latter determines the number of sentences to group together when evaluating semantic similarity. We set the breakpoint percentile threshold to 90, and the buffer size to 3.
@@ IV-B Results and Analysis
TABLE I: Accuracy comparison of our fine-tuned Phi-2 model against baseline models, both with and without retrieved context.
@@
In Table I, we benchmark our developed framework against three other solutions: base Phi-2 (2.7B), GPT-4o mini (8B), and GPT4o (1.76T). As expected, the base Phi-2 model shows poor performance of about 49.95% accuracy when tested on the dataset when no context is provided. Although the performance is notably improved to 71.35% when the base model is supplemented by retrieved context, it is still outperformed by our proposed model. Our fine-tuned RAG model is better aligned with the required task, enabling it to leverage the retrieved context better and produce more accurate recommendations.
Notably, our developed system also outperforms GPT-4o with and without context. Since GPT-4o is a significantly more generalized model, it is not well-aligned to the specialized task at hand. Qualitative analysis reveals that GPT-4o’s a priori knowledge from other domains sometimes interferes with our domain-specific task, albeit when the relevant context is present in the prompt.

### v2

@@ Abstract
Recent studies show that large language models (LLMs) struggle with technical standards in telecommunications. We propose a fine-tuned retrieval-augmented generation (RAG) system based on the Phi-2 small language model (SLM) to serve as an oracle for communication networks. Our developed system leverages forward-looking semantic chunking to adaptively determine parsing breakpoints based on embedding similarity, enabling effective processing of diverse document formats. To handle the challenge of multiple similar contexts in technical standards, we employ a re-ranking algorithm to prioritize the most relevant retrieved chunks. Recognizing the limitations of Phi-2’s small context window, we implemented a recent technique, namely SelfExtend, to expand the context window during inference, which not only boosts performance but also accommodates a wider range of user queries and design requirements from customers to specialized technicians. For fine-tuning, we utilize the low-rank adaptation (LoRA) technique to enhance computational efficiency during training and enable effective fine-tuning on small datasets. Our comprehensive experiments demonstrate substantial improvements over existing question-answering approaches in the telecom domain, achieving performance that exceeds larger language models such as GPT-4 (which is about 880 times larger in size). This work presents a novel approach to leveraging SLMs for communication networks, offering a balance of efficiency and performance. This work can serve as a foundation towards agentic language models for networks.
@@
**footnotetext: These authors contributed equally to this work.
@@ I Introduction
Recent studies indicate that while state-of-the-art LLMs perform well on general telecommunications queries, they struggle with questions related to technical standards in the field [11, 12]. We believe this is mainly due to the fact that standard-type knowledge and system specifications do not exist in common research papers and other publications, which serve as the main learning sources for LLMs. For instance, with respect to telecom systems, the polysemy of abbreviations (e.g., SAP: service access point vs. session announcement protocol) can hinder the model from inferring a correct answer. Additionally, the LMs training on generalized knowledge can interfere with its performance in specialized domain tasks. For instance, certain protocols and methods in telecom-specific domains do not necessarily follow the generally followed-upon practices in broader contexts. A model trained on generalized knowledge may not adequately capture these domain-specific nuances and practices.
We propose a carefully developed Phi-2-based fine-tuned RAG system to serve as an oracle for communication networks. To the best of our knowledge, this is the first work to present a fine-tuned RAG system for communication networks. Previous works focus on presenting a frozen RAG framework or fine-tuning an LM. Our RAG system leverages a forward-looking semantic chunking (or parsing) strategy that adaptively determines breakpoints between sentences based on embedding similarity. This approach enables the system to effectively process documents with diverse formatting. In the Third Generation Partnership Project (3GPP) documents, a query can often relate to multiple similar contexts, as discussions and paragraphs on related topics may appear in various sections or be phrased similarly. Therefore, we utilize a re-ranking algorithm to further rank the retrieved chunks based on their relevance to the input query.
@@ II Related Works
In [11], the introduction of the TeleQnA dataset marks a significant advancement in evaluating QnA tasks for telecommunications. The TeleQnA dataset contains 10,000 MCQs about telecommunication systems, curated from various sources such as 3GPP and research papers. The dataset is verified by telecom human-in-the-loop experts. Another dataset, namely TSpec-LLM, is recently released [16]. The authors develop an automated framework to generate QnA pairs from 3GPP specifications, then test a naive RAG architecture to assess the quality of their dataset. In [17], Gajjar et al. introduce ORAN-Bench-13K, a dataset dedicated to the evaluation of open radio access networks (O-RAN) tasks. The dataset is based on 116 O-RAN specification documents and contains 13,000 pairs of MCQs, based on which ORANSight is developed, a RAG-based framework.
@@ III-E Extending the Context Window with SelfExtend
[…]
SLM typically struggle to generalize effectively to input sequences longer than those encountered during training. This presents as a challenge during inference with long contexts. The context window of SLMs are often short. For example, Phi-2 has a context window of 2048 tokens. Semantic chunking does not guarantee a fixed chunk size, and therefore, might exceed the context window of the SLMs.
@@ III-F The Generator: Fine-tuned Phi-2 with Multiple Contexts
A substantial portion of telecom key terms and special language are confined to specification documents and white papers, neither of which LMs are heavily trained on. We fine-tune Phi-2 to overcome this and to adapt the model to recognize telecom terminology.
@@
The concept underlying LoRA is that pre-trained LMs possess a low ‘intrinsic dimension’ [26]. That is, the model’s essential information is concentrated in a smaller subspace, even if the overall parameter space is high-dimensional. LoRA harnesses this observation and focus the model updates on a smaller only a subset of the learning parameters as illustrated in Fig. 4. During fine-tuning, weight updates $W_{new}$ are represented as
@@ III-G Prompt Engineering
Prompt engineering is crucial for optimizing the performance of SLMs like Phi-2 in telecom applications. In this work, we supplement the inputted question with chunks of relevant context, and a set of instructions for the SLMs to follow, forming the prompt. An exemplary prompt is outlined in 3. This approach helps unify the format of the SLM’s output, focus its attention on the appearance of critical terms, and encourage it to reply on the context rather than its prior knowledge. This prompt results in focused outputs, ensuring accurate and relevant responses for complex telecom inquiries.
@@ IV-A Settings
The RAG framework relies on around 550 3GPP documents up to Release 18. For fine-tuning, we utilize the TeleQnA dataset which contains 10,000 multiple-choice questions [11]. The testing set features another 2,000 multiple-choice questions that focus on 3GPP standards. For fine-tuning, we set the following parameters: weight decay of 0.01, batch size of 32, dropout rate of 0.05, and learning rate of $10^{-4}$. For LoRA, we set the rank to 32 and the alpha to 64. The fine-tuning tagets the following layers: ’q_proj’,’k_proj’,’v_proj’, and ’dense’. For semantic chunking, there are two hyper-parameters, namely the breakpoint percentile threshold and the buffer size. The former represents the percentile of cosine dissimilarity that must be exceeded between a group of sentences and the next to form a node. The latter determines the number of sentences to group together when evaluating semantic similarity. We set the breakpoint percentile threshold to 90, and the buffer size to 3.
@@ IV-B Results and Analysis
TABLE I: Accuracy comparison of ours fine-tuned Phi-2 model against baseline models, both with and without retrieved context.
@@
In Table I, we benchmark our developed framework against three other solutions: base Phi-2 (2.7B) and GPT4o, a state-of-the-art model with extensive capacity. As expected, the base Phi-2 model shows poor performance of about 49.95% accuracy when tested on the dataset when no context is provided. Although the performance is notably improved to 71.35% when the base model is supplemented by retrieved-context, it is still outperformed by our proposed model. Our fine-tuned RAG model is better aligned with the required task, enabling it to leverage the retrieved context better and produce more accurate recommendations.
Notably, our developed framework performs on par with GPT-4o with and without context. Interestingly, the size of GPT-4o does not directly correlate with its performance on this specialized task. One possible explanation for this is that GPT-4o, being a highly generalized model, lacks precise alignment with the specific task requirements. Qualitative analysis suggests that GPT-4o’s a priori knowledge from other domains sometimes interferes with our domain-specific task, albeit when the relevant context is present in the prompt.